OpenStudy (anonymous):

What are the approximate solutions of 2x^2 + 5x = 4 rounded to the nearest hundredth?

OpenStudy (nathan917):

Simplifying 2x2 + 5x = 4 Reorder the terms: 5x + 2x2 = 4 Solving 5x + 2x2 = 4 Solving for variable 'x'. Reorder the terms: -4 + 5x + 2x2 = 4 + -4 Combine like terms: 4 + -4 = 0 -4 + 5x + 2x2 = 0 Begin completing the square. Divide all terms by 2 the coefficient of the squared term: Divide each side by '2'. -2 + 2.5x + x2 = 0 Move the constant term to the right: Add '2' to each side of the equation. -2 + 2.5x + 2 + x2 = 0 + 2 Reorder the terms: -2 + 2 + 2.5x + x2 = 0 + 2 Combine like terms: -2 + 2 = 0 0 + 2.5x + x2 = 0 + 2 2.5x + x2 = 0 + 2 Combine like terms: 0 + 2 = 2 2.5x + x2 = 2 The x term is 2.5x. Take half its coefficient (1.25). Square it (1.5625) and add it to both sides. Add '1.5625' to each side of the equation. 2.5x + 1.5625 + x2 = 2 + 1.5625 Reorder the terms: 1.5625 + 2.5x + x2 = 2 + 1.5625 Combine like terms: 2 + 1.5625 = 3.5625 1.5625 + 2.5x + x2 = 3.5625 Factor a perfect square on the left side: (x + 1.25)(x + 1.25) = 3.5625 Calculate the square root of the right side: 1.887458609 Break this problem into two subproblems by setting (x + 1.25) equal to 1.887458609 and -1.887458609. Subproblem 1 x + 1.25 = 1.887458609 Simplifying x + 1.25 = 1.887458609 Reorder the terms: 1.25 + x = 1.887458609 Solving 1.25 + x = 1.887458609 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-1.25' to each side of the equation. 1.25 + -1.25 + x = 1.887458609 + -1.25 Combine like terms: 1.25 + -1.25 = 0.00 0.00 + x = 1.887458609 + -1.25 x = 1.887458609 + -1.25 Combine like terms: 1.887458609 + -1.25 = 0.637458609 x = 0.637458609 Simplifying x = 0.637458609 Subproblem 2 x + 1.25 = -1.887458609 Simplifying x + 1.25 = -1.887458609 Reorder the terms: 1.25 + x = -1.887458609 Solving 1.25 + x = -1.887458609 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-1.25' to each side of the equation. 1.25 + -1.25 + x = -1.887458609 + -1.25 Combine like terms: 1.25 + -1.25 = 0.00 0.00 + x = -1.887458609 + -1.25 x = -1.887458609 + -1.25 Combine like terms: -1.887458609 + -1.25 = -3.137458609 x = -3.137458609 Simplifying x = -3.137458609 Solution The solution to the problem is based on the solutions from the subproblems. x = {0.637458609, -3.137458609}

OpenStudy (nathan917):

Let me know if You want me to round them :P

OpenStudy (anonymous):

No real solutions x ≈ −6.27 and x ≈ 1.27 x ≈ −0.64 and x ≈ 3.14 x ≈ −3.14 and x ≈ 0.64

OpenStudy (anonymous):

yea i need them rounded

OpenStudy (nathan917):

Its C.

OpenStudy (anonymous):

thnx

OpenStudy (campbell_st):

why not just use the general quadratic formula \[x = \frac{-5 \pm \sqrt{5^2 - 4 \times2\times -4}}{2 \times 2}\] just seems simpler to me...